OpenStudy (anonymous):

systems of equation: a-b=-2 -2a+b=7 (okay i get that we have to have a elimination method to solve it but how do we start it? if you know what i mean!)

OpenStudy (lgbasallote):

just add the two equations...b gets canceled out -a = 9 a = -9 b = -7

OpenStudy (anonymous):

so a-b=-2 9+(-7)+(-2) -2a+b=7 -2(9)+(-7)=2 ?????????

OpenStudy (lgbasallote):

^you misplaced the negative signs

OpenStudy (anonymous):

where???

OpenStudy (lgbasallote):

hahahha :)) crazygirl2012....i'm very sorry but the answer I gave was wrong...i was substituting the values to the second equation and it wasn't becoming equal to 7 the correct answer is a = -5 b = -3 I am really really really sorry. Checking: (-5) - (-3) = -2 -5 + 3 = -2 -2 = -2 -2(-5) + (-3) = 7 10 -3 = 7 7 = 7 so you know your values are correct.

OpenStudy (anonymous):

Thanks